OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve 3-5(x+2) = 3(x+1)+6 In your own words please include your work at each step.

zepdrix (zepdrix):

What part are you having trouble with Fatboi? Do you understand how to solve the equation or no? Or is it just explaining the steps that's confusing?

OpenStudy (anonymous):

What are you working on?

OpenStudy (anonymous):

Explaining the problem!

OpenStudy (anonymous):

First I would work with the brackets first.

zepdrix (zepdrix):

So a good first step would be to use the `Distributive Property` on each side. In your own words, you might say "I start by distributing the -5 to each term in the brackets." And so the same on the other side.

OpenStudy (anonymous):

Okay!

zepdrix (zepdrix):

\[\large 3-5(x+2) = 3(x+1)+6\]Giving us something like this, yes?\[\large 3-5x-10 = 3x+3+6\] You might want to be a little bit more thorough in there, explaining that you're multiplying the -5 and 2... I dunno :)

zepdrix (zepdrix):

It's suppose to be in `your words`, so it's hard to give you the right things to say hehe

OpenStudy (anonymous):

Right...

OpenStudy (anonymous):

That's what I have. I will figure out the right things to say! Thanks!

zepdrix (zepdrix):

cool c:

OpenStudy (anonymous):

You're so welcome!